# Training towards significance with the decorrelated event classifier transformer neural network

### v2

@@ IV.4 Machine learning technique evaluation metrics
    The search region is binned according to the machine learning technique output.
@@
    Normalized $m_{\ell\ell\gamma}$ histograms are created for each machine learning technique bin.
@@
    The differences between the machine learning technique bins is measured with the standard deviation of the normalized yield for each mass bin.
@@
When calculating the correlation, the sample is divided into 8 machine learning technique bins with equal number of signal events in each bin.
@@ IV.5 Experiment and results
A repeated random subsampling validation procedure [37] is used to evaluate the machine learning techniques, where 3 trials are done. The number of trials is limited due to the long training time of the neural networks. The validation procedure follows the steps below,
@@
    The events in the data set are randomly shuffled and divided equally for the training, validation, and test data set.
@@
    The machine learning technique is trained using the training data set. For neural networks, the best model between the epochs is selected using the expected significance metric on the validation data set.
@@
    The performance of the machine learning technique using the evaluation metrics is evaluated on the test data set.
@@
    The steps above are repeated $N$ times, where in this work $N=3$. The machine learning technique is re-initialized for each trial.
@@
    After all the trials, the evaluation metrics are averaged to assess the performance of the machine learning techniques.
The average evaluation metrics for the machine learning techniques are shown in Table 1. The experiment results show the following:
  • The event classifier transformer trained with DisCo loss shows the highest significance with lowest background correlation between the machine learning techniques.
@@
Table 1: Average evaluation metrics for machine learning techniques using the random subsampling evaluation procedure with 3 trials. Random is a random classifier. FNN is feed-forward network. DFNN is deep feed-forward network. ETN is event classifier transformer network. N.A. is not available. Ext is extreme loss. Signi. is expected significance. AUC is area under the curve of the receiver operating characteristic curve. Bkg. Corr. is the correlation of $m_{\ell\ell\gamma}$ with the machine learning technique output for the background calculated with the method described in section IV.4. Best epoch is the epoch that had the highest significance on the validation data set. Bold numbers indicate the best values over the machine learning techniques.
@@
There is an upper limit on performance for a classifier due to the similarities of the signal and background, where machine learning classifiers could be optimized in approaching the upper limit in a certain phase spaces. Therefore the goal for the machine learning classifier is to approach the upper limit in the phase space that is most relevant for the problem. The event classifier transformer tends to be able to approach the upper limit on the significance metric better than the other machine learning techniques used in this paper. However with more hypertraining on the other machine learning techniques or by using different machine learning techniques, it could be possible to get closer to the upper limit, where this is left for future studies.
@@
The machine learning technique metrics are shown for one of the subsampling trials. The correlation between the $m_{\ell\ell\gamma}$ and network output is shown in Fig. 5 for the deep feed-forward network trained with BCE loss and the event classifier transformer trained with the extreme + DisCo loss. The network output distribution is shown in Fig. 6, where in the figure overtraining of the network is measured with the $\chi^{2}$ test [38] implemented in ROOT [39], by comparing the network output distributions on the training sample and validation sample. The residuals in the plot are the normalized differences between the output distributions defined in [38]. For both networks, the bump in the middle of the signal distribution is related with the networks making different output distributions depending on $p_{T}^{\ell\ell\gamma}/m_{\ell\ell\gamma}$. An effect from the DisCo term can be seen, where the minimum network output value is shifted upwards.
@@
Figure 5: $m_{\ell\ell\gamma}$ distribution of the background, where each histogram is a bin in the machine learning technique output distribution. Each bin has an equal number of signal events. Lower signal percentile (sig. p) values correspond to higher network output values. (Above) Deep feed-forward network trained with BCE loss. (Below) Event classifier transformer network trained with extreme + DisCo loss. Correlation represents the magnitude of difference in the shapes between the machine learning bins. A lower correlation can be observed with the network trained with DisCo loss.
@@
Figure 6: (Left) Deep feed-forward network trained with BCE loss. (Right) Event classifier transformer network trained with the extreme + DisCo loss. (Top) Network output distribution on the training data set and validation data set. (Bottom) Overtraining of the network evaluated by comparing the network output distribution between the training and validation data set for the signal and background using a $\chi^{2}$ test, where residuals are the normalized differences between the output distributions.
@@
It is noted that when increasing the number of input features, the increase in the number of trainable network weights for the event classifier transformer is small compared to a deep feed-forward network. The smaller increase can make the network easier to train, when more input features are used. For example, when changing the number of input features from 12 features to 15 features, the increase of the number of trainable weights for the event classifier transformer is 265 (5% increase), while for the deep feed-forward neural network it is 4,671 (55% increase).
@@
A study was also performed to compare the results when the data scope training method was not applied. Instead, the wide mass window was used for both the classifier loss and DisCo loss. The training became difficult where the correlation could be high or the network outputs could converged to one value. The networks that were successfully trained had a few percent poorer significance, when having similar background correlations.
Figure 7: BCE loss (left) and expected significance (right) over training epochs for a feed-forward neural network trained with BCE loss. The epoch that has the lowest BCE loss does not correspond to the epoch that has the highest significance. The network has a high significance in the beginning of the training, and becomes lower as the training continues. This behavior is due to the neural network becoming more discriminative for neural network outputs at medium values but as a trade off becomes less discriminative for neural network outputs at high values, where the significance depends heavily on.
@@ V Related work
The Particle Transformer [8] is a transformer-based neural network targeted towards jet-flavor tagging. The structure is similar to the transformer [7], where each particle in a jet is considered as a token. Additionally, variables calculated using the features of a pair of particles are passed through a feed-forward network and added to the attention scores of the MHA. For the final prediction, class tokens are passed as a query to an MHA. The event classifier transformer in this work has a similar structure where the main difference is that each event feature is embedded to a separate token using separate feed-forward networks and that the neural network output is used to bin a bump hunt analysis.
There has been work to develop a loss function targeted towards significance, where a loss based on the inverse of the significance [40] has been studied. However when a network is trained with this loss, the neural networks outputs are clustered around 0 or 1. This behavior makes the loss unusable for binning a bump hunt analysis. The proposed extreme loss can enhance the significance, while not having the network outputs clustered around 0 or 1. It can also reduces the number of epochs that are needed to reach the optimal performance of the network.
DisCo [15] has been used for jet flavor tagging and for the ABCD analysis technique [41], where it has been shown to be effective in decorrelating a targeted feature of the jet with the neural networks output and decorrelating two neural network outputs used for the ABCD method. In this work, DisCo is used to decorrelate the neural network output with the reconstructed mass to bin a bump hunting analysis.
@@ VI Summary and conclusions
A transformer-based neural network is proposed to increase the expected significance for a search of a resonance in a reconstructed mass distribution. The significance is enhanced by binning events using a network that discriminates signal and background events. To apply the transformer architecture for this task, each event feature is passed through a separate feed-forward network to create tokens for the transformer. This network is called the event classifier transformer.
@@
In the context of a simplified $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ search, the new event classifier transformer trained with the special techniques shows higher significance and lower mass correlation when compared with boosted decision trees and feed-forward networks. This result demonstrates the potential of the event classifier transformer and the specialized training techniques targeted towards binning a search for a resonance in the reconstructed mass distribution.

### v3

@@ IV.4 Machine-learning technique evaluation metrics
    The search region is binned according to the machine-learning technique output.
@@
    Normalized $m_{\ell\ell\gamma}$ histograms are created for each machine-learning technique bin.
@@
    The differences between the machine-learning technique bins are measured with the standard deviation of the normalized yield for each mass bin.
@@
When calculating the correlation, the sample is divided into eight machine-learning technique bins with an equal number of signal events in each bin.
@@ IV.5 Experiment and results
A repeated random subsampling validation procedure [37] is used to evaluate the machine-learning techniques, where three trials are done. The number of trials is limited due to the long training time of the neural networks. The validation procedure follows the steps below:
@@
    The events in the dataset are randomly shuffled and divided equally for the training, validation, and test datasets.
@@
    The machine-learning technique is trained using the training dataset. For neural networks, the best model between the epochs is selected using the expected significance metric on the validation dataset.
@@
    The performance of the machine-learning technique using the evaluation metrics is evaluated on the test dataset.
@@
    The steps above are repeated $N$ times, where in this work $N=3$. The machine-learning technique is reinitialized for each trial.
@@
    After all the trials, the evaluation metrics are averaged to assess the performance of the machine-learning techniques.
The average evaluation metrics for the machine-learning techniques are shown in Table 1. The experiment results show the following:
  • The event classifier transformer trained with DisCo loss shows the highest significance with the lowest background correlation between the machine-learning techniques.
@@
Table 1: Average evaluation metrics for machine-learning techniques using the random subsampling evaluation procedure with three trials. Random is a random classifier. FNN is a feed-forward network. DFNN is a deep feed-forward network. ETN is an event classifier transformer network. N.A. means not available. Ext is extreme loss. Signi. is the expected significance. AUC is the area under the curve of the receiver operating characteristic curve. Bkg. Corr. is the correlation of $m_{\ell\ell\gamma}$ with the machine-learning technique output for the background calculated with the method described in Sec. IV.4. Best epoch is the epoch that had the highest significance on the validation dataset. Bold numbers indicate the best values over the machine-learning techniques.
@@
There is an upper limit on performance for a classifier due to the similarities of the signal and background, where machine-learning classifiers could be optimized in approaching the upper limit in a certain phase spaces. Therefore the goal for the machine-learning classifier is to approach the upper limit in the phase space that is most relevant for the problem. The event classifier transformer tends to be able to approach the upper limit on the significance metric better than the other machine-learning techniques used in this paper. However with more hypertraining on the other machine-learning techniques or by using different machine-learning techniques, it could be possible to get closer to the upper limit, which is left for future studies.
@@
The machine-learning technique metrics are shown for one of the subsampling trials. The correlation between the $m_{\ell\ell\gamma}$ and network output is shown in Fig. 5 for the deep feed-forward network trained with BCE loss and the event classifier transformer trained with the extreme + DisCo loss. The network output distribution is shown in Fig. 6, where in the figure overtraining of the network is measured with the $\chi^{2}$ test [38] implemented in ROOT [39], by comparing the network output distributions on the training sample and validation sample. The residuals in the plot are the normalized differences between the output distributions defined in [38]. For both networks, the bump in the middle of the signal distribution is related with the networks making different output distributions depending on $p_{T}^{\ell\ell\gamma}/m_{\ell\ell\gamma}$. An effect from the DisCo term can be seen, where the minimum network output value is shifted upwards.
@@
Figure 5: $m_{\ell\ell\gamma}$ distribution of the background, where each histogram is a bin in the machine-learning technique output distribution. Each bin has an equal number of signal events. Lower signal percentile (sig. p) values correspond to higher network output values. Top: deep feed-forward network trained with BCE loss. Bottom: Event classifier transformer network trained with extreme + DisCo loss. Correlation represents the magnitude of difference in the shapes between the machine-learning bins. A lower correlation can be observed with the network trained with DisCo loss.
@@
Figure 6: Left: deep feed-forward network trained with BCE loss. Right: event classifier transformer network trained with the extreme + DisCo loss. Top: network output distribution on the training dataset and validation dataset. Bottom: overtraining of the network evaluated by comparing the network output distribution between the training and validation datasets for the signal and background using a $\chi^{2}$ test, where residuals are the normalized differences between the output distributions.
@@
It is noted that when increasing the number of input features, the increase in the number of trainable network weights for the event classifier transformer is small compared to a deep feed-forward network. The smaller increase can make the network easier to train, when more input features are used. For example, when changing the number of input features from 12 to 15, the increase of the number of trainable weights for the event classifier transformer is 265 (a 5% increase), while for the deep feed-forward neural network it is 4,671 (a 55% increase).
@@
A study was also performed to compare the results when the data scope training method was not applied. Instead, the wide mass window was used for both the classifier loss and the DisCo loss. The training became difficult where the correlation could be high or the network outputs could converge to one value. The networks that were successfully trained had a few-percent poorer significance when having similar background correlations.
Figure 7: BCE loss (left) and expected significance (right) over training epochs for a feed-forward neural network trained with BCE loss. The epoch that has the lowest BCE loss does not correspond to the epoch that has the highest significance. The network has a high significance in the beginning of the training, and it becomes lower as the training continues. This behavior is due to the neural network becoming more discriminative for neural network outputs at medium values, but as a tradeoff, it becomes less discriminative for neural network outputs at high values, where the significance dependence is high.
@@ V Related work
The Particle Transformer [8] is a transformer-based neural network targeted toward jet-flavor tagging. The structure is similar to the transformer [7], where each particle in a jet is considered as a token. Additionally, variables calculated using the features of a pair of particles are passed through a feed-forward network and added to the attention scores of the MHA. For the final prediction, class tokens are passed as a query to a MHA. The event classifier transformer in this work has a similar structure where the main difference is that each event feature is embedded to a separate token using separate feed-forward networks and that the neural network output is used to bin a bump hunt analysis.
There has been work to develop a loss function targeted toward significance, where a loss based on the inverse of the significance [40] has been studied. However when a network is trained with this loss, the neural network outputs are clustered around 0 or 1. This behavior makes the loss unusable for binning a bump hunt analysis. The proposed extreme loss can enhance the significance while not having the network outputs clustered around 0 or 1. It can also reduce the number of epochs that are needed to reach the optimal performance of the network.
DisCo [15] has been used for jet flavor tagging and for the ABCD analysis technique [41], where it has been shown to be effective in decorrelating a targeted feature of the jet with the neural network output and decorrelating two neural network outputs used for the ABCD method. In this work, DisCo is used to decorrelate the neural network output with the reconstructed mass to bin a bump hunting analysis.
@@ VI Summary and conclusions
A transformer-based neural network is proposed to increase the expected significance of a search for a resonance in a reconstructed mass distribution. The significance is enhanced by binning events using a network that discriminates between signal and background events. To apply the transformer architecture for this task, each event feature is passed through a separate feed-forward network to create tokens for the transformer. This network is called the event classifier transformer.
@@
In the context of a simplified $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ search, the new event classifier transformer trained with the special techniques shows higher significance and lower mass correlation when compared with boosted decision trees and feed-forward networks. This result demonstrates the potential of the event classifier transformer and the specialized training techniques targeted toward binning a search for a resonance in the reconstructed mass distribution.